# Airfoil Diffusion: Denoising Diffusion Model For Conditional Airfoil Generation

### v2

@@ 5 Conclusion
In this paper, we presented a novel diffusion-based framework for airfoil design, leveraging the UIUC airfoil database to train a model capable of generating a wide variety of airfoil profiles. By conditioning the generation process on specific aerodynamic coefficients such as lift ($C_{l}$) and drag ($C_{d}$), our model demonstrated the ability to produce airfoils with desirable aerodynamic and geometric criteria.
@@
One of the key strengths of our approach lies in its capacity to produce novel airfoil geometries, as evidenced by the Chamfer Distance analysis. The generated airfoils, while grounded in the characteristics of the training data, exhibit significant novelty, indicating the model’s potential for discovering innovative aerodynamic shapes. Importantly, the generated designs remain within a realistic and aerodynamically viable space, ensuring their practical applicability.
In summary, the diffusion model-based framework we introduced represents a powerful tool for expanding the design space and discovering new, high-performance airfoil shapes. Future work could explore extending this methodology to three-dimensional aerodynamic shapes, integrating multiple conditioning parameters within a single model, and applying this generative approach to other domains where innovative design solutions are critical.

### v3

@@ 5 Conclusion
In this paper, we presented a novel diffusion-based framework for airfoil design, leveraging the UIUC airfoil database to train a model capable of generating new designs within the training sample design space. By conditioning the generation process on specific aerodynamic coefficients such as lift ($C_{l}$) and drag ($C_{d}$), our model demonstrated the ability to produce airfoils with desirable aerodynamic and geometric criteria not seen in training samples.
@@
One of the key strengths of our approach lies in its capacity to produce novel airfoil geometries, as evidenced by the PCA reduction analysis. The generated airfoils, while grounded in the characteristics of the training data, exhibit moderate novelty, indicating the model’s potential for discovering innovative aerodynamic shapes. Importantly, the generated designs remain within a realistic and aerodynamically viable space, ensuring their practical applicability.
In summary, the diffusion model-based framework we introduced represents a powerful tool for exploring the design space and discovering new, high-performance airfoil shapes. Future work could explore extending this methodology to three-dimensional aerodynamic shapes, integrating multiple conditioning parameters within a single model, and applying this generative approach to other domains where innovative design solutions are critical.